OpenStudy (anonymous):

Write an expression that is equivalent to two over three (4x + 9). A. eight over threex + 6 B.eight over threex + 9 C.two over threex + 6 D.two over threex + 9

OpenStudy (anonymous):

PLEASE HELP

jagr2713 (jagr2713):

first we distribute \[\frac{ 2 }{ 3 }\times 4x+\frac{ 2 }{ 3 }\times 9\]

jagr2713 (jagr2713):

what do you think we o next @Hanaabdalah

OpenStudy (anonymous):

i didn't understand

jagr2713 (jagr2713):

ok we set it up like that so we distributed so now in that form what do we do with the 2 and 4

jagr2713 (jagr2713):

Dont we multiply them? @Hanaabdalah

jagr2713 (jagr2713):

@is3535 this is your second warning.It is better to guide the user to an answer rather than doing all the work yourself. We discourage giving out answers without letting the asker at least take some part in the analysis - the aim of this site is to teach. Please refer to our code of conduct at http://openstudy.com/code-of-conduct for more information. Thank you.

OpenStudy (is3535):

srry sir

jagr2713 (jagr2713):

Do you till want the explanation so you can do answers like this in the future or no @Hanaabdalah

OpenStudy (is3535):

do you need more help @Hanaabdalah

OpenStudy (anonymous):

yes plz